the sum of the least and greatest of three consecutive odd integer - 32
Answer by venugopalramana(3286)

take n as any number . 2n will be even . add one to it to get odd number.the next oddnumber will be 2 more than this.(EX.1,3,5...ETC).so let the 3 consecutive odd integers br 2n+1,2n+3,2n+5
sum of the least and greatest=2n+1+2n+5=4n+6=-32
4n=-32--6=-38
n=-38/4=no solution..there are no three consecutive integers satisfying the above condition.